National Income Accounts.
X 100. Transfers 100. PIT 220.
SST 60. IBT1 40. NPI 40. I 100.
CRE 20. C 700. NFFI 100. Depreciation 60.
IBT2 40. G 300. M 200.
Determine DP
1
Expert's answer
2016-05-16T12:38:03-0400
The DP is defined as follows:
DP =C+G+I+(X-M)
DP =700 + 300 + 100 + (100-200) = 1000